Integrate g(x, y) = x⁴ + y² over the region 

bounded by y = x, y = 2x and x = 2. 


1
Expert's answer
2021-05-04T12:43:35-0400

Using double integration with respect to dydx integrate x4 +y2 we start with dy

Integrating with respect to dy we get

x4y + y3/3 limit from x to 2x substituting the limit we get x5 +7x3/3

Then integrate with respect to dx we get x6/6 +7x4/12 substituting the limit 0-2 the solution becomes 64/6 +112/12= 240/12=20

The answer becomes 20